OpenStudy (andrewkaiser333):

Maggie recently took a road trip. She bought 15 gallons of gasoline for $1.68 per gallon and 12 gallons at a price of $1.76 per gallon. She filled her tank once on the way back with 16 gallons at $1.65 per gallon. What was the average fuel cost per gallon on Maggie's trip?

OpenStudy (andrewkaiser333):

@mathstudent55

OpenStudy (baru):

average fuel cost per gallon = (total cost) / (total number of gallons) total cost = (1.68 X 15) + (1.76*12) + (1.65*16) = $72.72 total no of gallons = 15+12+16 =43 average = 72.72 / 43 = $1.69